OpenStudy (wolfadaskies):

what is the hypotenuse of a triangle with the measurements of 14 and 11? 12 and 20? 26 and 20? Can the set of lengths be the side lengths of a right triangle? 8 ft, 15 ft, 17 ft

OpenStudy (anonymous):

do you know how to use the Pythagorean theorem?

OpenStudy (wolfadaskies):

Yes, a^2+b^2=c^2. But I think I'm just way overthinking this one, heh.

OpenStudy (wolfadaskies):

@KlOwNlOvE

OpenStudy (anonymous):

Which letter represents the hypotenuse

OpenStudy (anonymous):

\[14^{2}+11 ^{2}=c ^{2}\]

OpenStudy (anonymous):

That is an example on how to setup number one and you can work through the rest based off that, goodluck!

OpenStudy (wolfadaskies):

Got the first one; it's 17.8 - can you help me with the others?

OpenStudy (anonymous):

12^2+20^2=c^2